### v1

Abstract
Coherent elastic neutrino-nucleus scattering (CENS) provides a unique probe for neutrino properties Beyond the Standard Model (BSM) physics. REactor neutrino LIquid xenon Coherent Scattering experiment (RELICS), a proposed reactor neutrino program using liquid xenon time projection chamber (LXeTPC) technology, aims to investigate the CENS process of antineutrinos off xenon atomic nuclei. In this work, the design of the experiment is studied and optimized based on Monte Carlo (MC) simulations. To achieve a sufficiently low energy threshold for CENS detection, an ionization-only analysis channel will be adopted for RELICS. A high emission rate of delayed electrons after a big ionization signal is the major background, leading to an analysis threshold of 120 photo-electrons in the CENS search. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. The physics potential of RELICS was explored with a 32 exposure at a baseline of 25 m from a reactor core with a 3 GW thermal power. In an energy range of 120 to 240 PEs, we the expected 13673.5 CENS and 2133.7 background events.The sensitivity of RELICS to the weak mixing angle was investigated at a low momentum transfer. Our study has shown that RELICS can further improve the constraints on the non-standard neutrino interaction (NSI) compared to the best results. RELICS set out to develop a powerful surface-level detection technology for low-energy neutrinos from reactors.
- PACS numbers
- Keywords
pacs:
95.35.+d, 14.80.Ly, 29.40.-n, 95.55.VjI Introduction
Coherent elastic neutrino-nucleus scattering (CENS) is a Standard Model (SM) weak neutral current process first proposed in 1974 [1]. The neutrino treats the nucleus as a single particle during scattering with a cross-section proportional to the nucleon number squared. Despite this process having a relatively large cross-section, the detection is challenging due to the ultra-low deposited energy. This process had evaded experimental discoveries for decades until 2017 when it was first observed by COHERENT collaboration with neutrinos produced by the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory scattering off a CsI scintillator [2], and later confirmed with a liquid argon detector [3] in 2020. These measurements of CENS offer new probes for diverse physical phenomena.
CENS can provide new constraints on physics beyond the Standard Model (SM). The weak mixing angle is related to the CENS cross-section [4]. Any deviation from the SM prediction in measurement can be an indication of new physics. CENS is also a valuable tool to constrain nonstandard neutrino interactions (NSI) [5], to search for sterile neutrinos [6], to verify the existence of non-trivial neutrino electromagnetic properties [7], and to study the nuclear structure [8] and astrophysics [9]. In addition, the CENS signal from the solar or atmospheric neutrinos is expected to be an irreducible background for the next-generation dark matter direct detection experiments, such as PandaX-30T [10] and XLZD [11]. A precise measurement of this process can provide valuable information for dark matter experiments digging into the so-called “neutrino floor/fog” [12, 13].
The observation of CENS by COHERENT collaboration inspired the interest in CENS experimental study. While the neutrino beam has been the primary source for the CENS discovery, it is limited by the total flux delivered and the availability of the facility. Therefore, nuclear reactors, with large anti-neutrino fluxes delivered with typical energies of a few MeV, which is the so-called “fully coherent regime” of CENS , have been proposed as attractive alternative facilities for the CENS study. Due to the considerably low recoil energy and complex background, the reactor CENS is harder to detect. Many efforts are underway with reactor neutrinos, such as [14, 15, 16, 17, 18, 19, 20, 21, 22, 23]. Detecting reactor CENS will provide a wealth of physics opportunities and potential applications such as nuclear safeguards [24].
RELICS is a reactor neutrino experiment using the LXeTPC technology. LXeTPC is one of the leading technologies for rare event searches, such as detecting dark matter and neutrinoless double beta decay [11, 25, 26, 27, 28, 29]. Benefiting from its ultra-low threshold by using ionization-only channel [30, 31] and reduced background, LXeTPC is an excellent approach for reactor CENS detection. The RELICS experiment is proposed to be located near the Sanmen Nuclear Power Plant in Zhejiang province, China. At a 25 m baseline to the 3 GW reactor core, the average neutrino flux available for RELICS is 1013cm-2s-1.
This work is organized as follows. The operation principle and the design of the detector as well as the shielding system, are described in detail in section II. The simulation framework is introduced in section III. In section IV, we predict the expected CENS event rates from the Sanmen nuclear reactor. Section V discusses a dedicated simulation of background contributions. The physics potential of RELICS is provided in section VI. A summary is presented in section VII.
II experimental designs
II.1 RELICS detector
The LXeTPC technology measures the energy deposited from particle interactions by simultaneously detecting the prompt scintillation light (S1) and delayed ionization charge (S2) [32]. The S1 is produced by the direct excitation of xenon atoms and the recombination of electron-ion pairs in LXe. The S2 is proportional to the number of liberated electrons extracted into gas xenon (GXe), which are then amplified via electroluminescence. Both S1 and S2 are recorded by photosensors, usually photomultiplier tubes (PMTs), on top and bottom arrays of the TPC. This allows a three-dimensional (3D) vertex reconstruction with a sub-centimeter precision, and the S2/S1 ratio can be used to discriminate between nuclear recoil (NR) and electronic recoil (ER) for background reduction. Typically, the energy threshold of LXeTPC is limited by the S1 signal in 23 photo-electrons (PEs). The inherent S2 amplification provides an alternative approach to decrease the energy threshold by utilizing the ionization-only channel, usually called the “S2-only” analysis. This approach can reduce the energy threshold to a few ionization electrons (equivalent NR energy below 1 keVnr) by sacrificing the NR/ER discrimination and depth information, which, in the case of CENS detection, can be largely compensated by the high signal-to-background ratio profited from the intense reactor neutrino flux.
A preliminary 3D mechanical design of the RELICS detector is shown in figure 1. A cylindrical TPC of 24 cm in height and 28 cm in diameter sits in a double-layer stainless steel (S.S) vessel, which provides a stable cryogenic environment. The target volume containing an active LXe of approximately 44 kg is defined by 12 interlocking and light-tight PTFE (polytetrafluoroethylene, Teflon) panels. The active LXe is then viewed by two arrays of 64 Hamamatsu R8520-406 PMTs on top and bottom. A fiducial mass of 32 kg LXe with only a radius selection of 12 cm can be well-defined for CENS search under “S2-only” analysis. A “diving bell” structure is mounted above the top PMT array to control the liquid level. Five etched S.S electrodes, including two grounding screenings on top and bottom to protect PMTs from high voltage (HV), anode, gate, and cathode to provide a nominal drift field of 500 V/cm and an extraction field of 10 kV/cm in GXe, are installed. To ensure the uniformity of the drift field, the target volume is encircled by several field-shaping rings made from oxygen-free high thermal conductivity (OFHC) copper. The TPC is surrounded by approximately 62 kg of LXe, which acts as a 4 active veto for background suppression. The LXe veto volume is instrumented with 48 R8520-406 PMTs. To achieve a high light collection, the inner surface of the inner vessel is covered with a thin PTFE foil. It is worth noting that the LXe veto is designed to reduce the background by 5 times by removing the coincidence events inside TPC and in the LXe veto layer. All materials for the detector construction will be carefully selected to ensure low intrinsic radioactivity.
II.2 Shielding system
Near-reactor neutrino experiments usually operate with shallow overburdens. Thus, the high-energy cosmic-ray neutrons are supposed to be one of the main backgrounds for the NR searches from CENS. A sufficient neutron shielding structure, such as water, is required to slow down and absorb the neutrons. Meanwhile, the cosmic muon and its induced secondaries are expected to be another major background. A highly efficient active muon veto anti-coincidence system is also critical to the success of this type of experiment.
Figure 2 shows the design of the RELICS shielding system based on Monte Carlo (MC) simulations, which is detailed in section V. The design has been optimized to satisfy the requirement of the CENS detection. The LXe detector is placed inside a shield of 7 m7 m7 m of water contained in a S.S tank. A 5 m water shield on top can suppress the cosmic-ray neutron-induced background to a controlled level. With such a water tank, the external neutrons and gammas from the laboratory environment and the nuclear reactor can be reduced to negligible levels. In close to the LXe detector, a muon veto detector is installed. The veto detector is made of 8 cm thick plastic scintillator (PS) equipped with wavelength-shifting fibers and silicon photomultipliers (SiPMs) to collect and detect the scintillation light. A high veto efficiency of 99% is expected for the RELICS experiment.
III Simulation framework
The MC simulation framework of the RELICS experiment (RelicsSim) is built upon a Geant4-based BambooMC toolkit [33]. Full detector geometry is implemented, including the detailed detector components and the shielding system. Various types of particle interactions are studied by simulation to evaluate the expected signal and background. Subsequently, RelicsAPT, which inherits from the Appletree [34] framework developed by XENONnT collaboration [35], is adopted to convert the energy depositions into observable S1 and S2 signals. The conversion is based on the intrinsic light yield (Ly) and charge yield (Qy) in LXe, and convoluted with the detector-related signal detection and reconstruction effects.
III.1 Liquid xenon response
Particle interaction in LXe will produce photons and electrons. The NEST model can characterize this process [36] after considering the fluctuations in scintillation, ionization, electron-ion recombination, and their drift field dependence. In RelicsAPT, the Ly and Qy from NEST (v2.3.6) at 500 V/cm drift field are used with no uncertainty assigned for NRs larger than 3 keVnr. To quantify the uncertainty of light and charge yields for CENS events, the Ly and Qy bands in XENON1T solar neutrino search [37] are adopted for NRs less than 3 keVnr. The bands of Ly and Qy are scaled to make the center curve connected to NEST at 3 keVnr. The final Ly and Qy of NRs applied in RelicsAPT are shown in figure 3. For ERs, Ly and Qy from NEST (v2.3.6) are used without uncertainties. In the 4 LXe veto region, both Ly and Qy are modeled with zero electric fields.
III.2 Detector related effects
In addition to the intrinsic LXe response model, the detector-related effects must be considered when interpreting the observed S1 and S2 signals. The main detector parameters are considered based on the expected detector performance listed in Tab. 1, as a conservative estimation.
To explore the capability of the RELICS detector on the ionization-only detection channel, the light collection efficiency (LCE) of S2 is simulated by RelicsSim with the optical parameters listed in Tab. 1. The LCE map is shown in figure 4. An averaged value of 30 PE/e- is estimated based on the S2 LCE map, the properties of the PMTs, such as , , and that also listed in Tab. 1, and a 100% electron extraction efficiency on the liquid surface. An electron lifetime of 1.2 s is applied for S2 estimation considering the probability of electrons loss to electronegative impurities during their drift.
| Optical parameters | |
| PTFE reflectivity (LXe & GXe) | 99% [38] |
| LXe Rayleigh scattering length | 30 cm [39] |
| LXe absorption length | 50 m [40] |
| Electrodes reflectivity | 57% [41] |
| Signal generation | |
| PMT quantum efficiency () | 33% [42] |
| PMT collection efficiency () | 75% [42] |
| Double PE probability () | 21.9% [43, 44] |
| 30 PE/e- | |
| Detector operation | |
| Drift field | 500 V/cm |
| Electron lifetime | 1.2 s [25] |
IV Expected CENS events
The SM of particle physics well predicts the CENS process — a neutrino of any flavor scatters off a nucleus via boson exchange coherently over the whole nucleus at a low momentum transfer (). The differential cross-section can be written as Eq. 1 below:
where is Fermi’s coupling constant, is the weak nuclear charge, is the mass of the target nucleus, is the ground state elastic form factor, is the nuclear recoil energy while the incident neutrino energy. is required to ensure coherence, where is the nuclear radius. Normally, this can be satisfied when MeV.
For neutrinos scattering in a target medium, the expected CENS event rate can be expressed as Eq. 2, where is the number of targets per unit mass, is the incident neutrino flux.
For the reactor neutrinos with energy above 2 MeV, the Huber-Mueller model[45, 46] is adopted for the spectra calculation while P. Vogel’s theory [47] is used for below 2 MeV region, which is rarely searched due to the 1.8 MeV energy threshold of inverse beta decay reaction. We assume a fuel mixture of 235U (56.1%), 238U (7.6%), 239Pu (30.7%), and 241Pu (5.6%), the total neutrino flux can be estimated with a 3 GW reactor core at 25m distance, and then the CENS event rate can be obtained accordingly as shown in figure 5.
V Background contributions
The complexity of the backgrounds in reactor CENS detection is one of the major reasons it escaped the searches for several decades. To overcome this challenge, dedicated background simulations have been conducted to optimize the design of the shielding system for RELICS, using RelicsSim and RelicsAPT. The background contributions can be divided into five categories according to their origin, such as (V.1) cosmic-ray neutrons, (V.2) cosmic muons, (V.3) detector materials, and (V.4) internal background from LXe. The last one is (V.5) delayed electrons, which have been observed by large LXeTPC searching for dark matter and should be treated carefully.
This section describes all the background components and is summarized in (V.6). The following criteria are applied to MC data to select the background events that can mimic CENS signals:
i.
LXe-veto: The 4-LXe veto exhibits a good background suppression because the background events are more likely to deposit energy in both active LXe and 4-LXe veto regions. The event will be removed if it contains an NR (ER) energy larger than 500 (100) keV in the LXe veto region.
ii.
single-scatter: Given the small interaction cross-section of neutrino, the probability of a neutrino having more than one interaction in the sensitive LXe volume is negligible. We require only one energy deposition above 0.5 keVnr energy in the LXe volume.
iii.
FV: Events are required to be within the center 12 cm radius fiducial to reduce background at the edge of the TPC, leading to a total fiducial volume (FV) of 32 kg.
iv.
S2-width: For “S2-only” analysis, the event depth cannot be accurately estimated. S2-width cut was developed to remove the events around the liquid-gas interface, especially for the backgrounds from detector materials.
v.
In addition, dedicated selection methods are developed to suppress backgrounds out of the delayed electrons, such as waveform classifier and pattern classifier. These two criteria are described in detail in (V.5).
V.1 Background induced by cosmic-ray neutrons
The cosmic rays that travel through the atmosphere can produce a variety of radiation particles, such as neutrons, protons, pions, etc, through spallation reaction on nitrogen, oxygen, etc [48, 49]. Among them, the high-energy cosmic-ray neutrons (CRNs) can easily pass through the shield and deposit energy in the RELICS detector. The main purpose of the water shield design is to block the CRNs. The size of the water tank is optimized based on the MC simulation. The CRN flux is calculated using the Cosmic-Ray Shower Generator (CRY) source [50]. The energy spectrum of the CRNs from CRY is shown in figure 6. Analyze the MC data shows that only CRNs with energy greater than 103 MeV can deposit energy in LXe, which also indicates that the neutrons from the nuclear reactor and laboratory environment normally with energy less than 10 MeV [51] that also shows in figure 6, can not contribute to the backgrounds.
After processing the MC data, and passing the LXe-veto, single-scatter and FV selections, the CRNs-induced NR background rate is estimated to be (7.7 0.7) in keVnr region-of-interest (ROI) of CENS detection, with the spectrum shows in figure 7 (top). In addition, the secondaries of CRNs will also induce ER background in LXe, with a rate of (9.0 0.3), which is shown in figure 7 (bottom).
V.2 Background induced by cosmic muons
The RELICS detector will operate without overburden. Hence, special attention has to be paid to the muon-related background. Even though muons are tagged with 99% efficiency by the PS veto detector, the neutron background produced by muon interacting with the detector materials is still one of the main background components. Dedicated simulations are performed to evaluate the muon-induced backgrounds. The energy and angular distributions of muons follow the Shukla model [52]. The energy deposition in LXe that is caused by muon-induced fast neutron and related gammas are recorded by RelicsSim, and then filtered with LXe-veto, single-scatter and FV selections. Assuming a 99% muon veto tagging efficiency, the muon-induced NRs and ERs are (0.08 0.01) in keVnr and (0.66 0.09) in keVer, respectively. The corresponding spectra are shown in figure 7.
V.3 Neutrons and gammas from the detector materials
Materials constructing the detector contain radioactive contamination. The selection of the materials requires an extensive radioactivity screening campaign for rare event search experiments. To evaluate the background contribution from these materials, the radioactive contamination for all the materials is considered in the RELICS MC model. Currently, the radioactivity of isotopes from XENON100 [53] is adopted. A dedicated effort is also being carried out to select clean materials for the RELICS experiment, leading to similar radioactivity for most of the materials.
The NR background rate induced by neutrons is around one event per year and completely negligible in the CENS ROI. ER background induced by gammas is studied in this work. The decay of these isotopes is generated and confined uniformly in the corresponding detector component in the simulation. By applying successive selection criteria, including LXe-veto, single-scatter and FV selections, the spatial distribution of ER background events inside the active LXe volume in the energy region of [0, 40] keVer is shown in figure 8. Figure 7 (bottom) shows the corresponding energy spectrum, which is two orders of magnitude higher than the ER background components. The average ER background rate is (364 5) in the FV volume.
S2 width selection: As shown in figure 8, a large amount of background events are located close to the liquid-gas interface. These backgrounds are mainly from the radioactivities of PMTs and can not be suppressed without the selection in “z” in an “S2-only” analysis. As demonstrated in [30], the pulse shape of S2 signals can discriminate background happening at the liquid-gas interface to signals distributed uniformly in “z”. A toy MC was conducted to consider the drift and diffusion properties of electrons in LXe to optimize such a selection. We simulate S2s and their pulse shape parameters with “z” distributions shown in figure 8 with energy in the CENS ROI. The S2s are simulated via RelicsAPT where the NEST model is incorporated. The S2 pulse shape is simulated using the electron diffusion model , where is the longitudinal diffusion coefficient, is the electron drift velocity, is the width of single electron waveform [54]. In our simulation. = 16 cm2/s, = 0.174 cm/s, and = 0.19s.
Uniformly distributed CENS signal is simulated using the same procedure. The distributions of the S2 width are shown in figure 9, for both CENS signal and ER background from the detector materials. A threshold of 0.22 on S2 width can give an average background rejection power of 94%, with a high CENS signal acceptance of 86%.
V.4 Internal backgrounds in LXe
Radioactivity in LXe can directly deposit energies and produce background in the CENS ROI. These radioactivities mostly come in the form of noble gases, such as argon, krypton, xenon, and radon. This section describes how these radioactivities are created and evaluates their contribution to the CENS ROI.
V.4.1 Intrinsic 222Rn and 85Kr
Krypton mixes into xenon gas when it is extracted from the air. 85Kr, regardless of its low abundance of in natural krypton, is one of the main backgrounds in any LXe experiments searching for rare events. The concentration of krypton in xenon is usually reduced below ppt level in dark matter search experiments, such as XENONnT and PandaX-4T, through distillations with dedicated columns[ref]. In RELICS, we assume a natKr/Xe concentration of 10 ppt and expect (0.28 0.01)10 as shown in figure 7 (bottom), which is a subdominant component.
The other intrinsic background in LXe comes from the decays of 222Rn, which emanates from any material touching Xenon gas or liquid. The emanated 222Rn diffuses homogeneously inside the LXe volume within its relatively long half-life of 3.8 days. The decay of 214Pb is another important background component in the CENS ROI in the decay chain of 222Rn. With similar size and material, the radioactivity of 222Rn in RELICS is comparable to XENON100. Assuming a conservative 222Rn radioactivity of 40 as achieved by the XENON100 experiment [53] (20 ), the expected ER background rate is estimated as of (0.98 0.02)10-3 kg-1day-1keV-1 as shown in figure 7 (bottom), which is also a subdominant component in the CENS search.
V.4.2 Cosmogenic isotopes and
The RELICS experiment, which is to be operated without overburden, is unavoidable under cosmic radiation. We consider cosmogenic radioactive isotopes through the activation and spallation of Xenon atoms. In particular, and are known cosmogenic backgrounds in LXeTPCs [55, 27] due to their characteristic x-rays or Auger electrons through electron capture process and their relatively long half-life of 36.4 and 35.0 days, respectively.
: Natural xenon does not contain . can be produced from when cosmogenic muons kick off a neutron in the nucleus with a cross-section of (2.74 0.4) cm2. Additionally, we consider another production channel through the neutron capture on with an upper limit cross-section of 0.7cm2 at 95% C.L [56]. The production rate of in LXe is calculated based on its production cross-section, the muon’s fluxes and energy spectrum, and thermal neutrons. The decays to via electron capture (EC) process, with a half-life of 36.4 days. We assume that the production and decay of reach equilibrium before deploying the RELICS detector inside the shielding. The estimated equilibrium decay rate is 9.29 evts/day/kg. The shielding system of RELICS can attenuate neutrons efficiently, and the equilibrium decay rate can decrease to 0.79 evts/day/kg after 300 days.
Following the EC, the is left primarily populating at the excited state of either 375 or 203 keV, which promptly decays to the ground state with gamma-ray emissions. The EC occurs from either the K, L, M, or N shell electrons and is subsequently filled with electrons from higher levels via emission of x-rays or Auger electrons with total energy (relative intensity) of 32.2 keV (83.4%), 5.2 keV (13.1%), 1.1 keV (2.9%) or 186 eV (0.7%), respectively. The final background contribution from is simulated with RelicsSim. The double-scatter criteria can remove most of these background events due to the simultaneous gamma-ray and x-ray or Auger electrons emission. If EC capture is happening from the N-shell electrons with a binding energy of 186 eV, it may contribute background in the CENS ROI when the emitted gamma rays, either with energy of 375 and 203 keV, escape from the LXe detector and the active LXe veto. The event rate from 186 eV energy deposition is less than kg-1day-1 after all the selection criteria described above, which is negligible for CENS search.
: We consider three possible mechanisms for production in xenon, following the approaches in [57]. Firstly, the extremely low content of in the atmosphere, which is generated by spallation of through (n, 4n), neutron capture of , or cosmic bombardment of calcium-containing soils, via (n, ). This mechanism only needs to be considered in the context of air leaks of an LXe detector, which is strictly limited to below 0.1 liter per year to ensure the purity level of LXe. Secondly, the spallation of detector materials, especially iron in steel. However, its contribution is greatly suppressed due to the slow diffusion rate of argon in steel. Lastly, the direct spallation of xenon atoms by high-energy cosmogenic protons and neutrons. The semi-empirical formula by Silberberg and Tsao was used to calculate the spallation cross-section. The CRY is used to calculate the proton spectrum. Gordon’s neutron spectrum [58] is adopted for cosmogenic neutron estimation. The production rate of this process is estimated to be 0.024 atoms/kg/day with an uncertainty of 100%. The produced decays to the ground state of via EC and subsequently deposits energy via x-rays or Auger electrons. The total energy (relative intensity) for K, L, and M shells are 2.82 keV (90.2%), 0.270 keV (8.9%), and 0.018 keV (0.9%), respectively. As a results, the background rate of in the CENS ROI is below kg-1day-1 and completely negligible for CENS search.
V.5 Delayed electrons
Besides the background components induced by real physical sources, another type of background must be considered in the CENS ROI. This background is caused by a high emission rate of delayed electrons after S2 signals in the TPC, which could be incorrectly identified as CENS candidate in “S2-only” analysis. Such background has been observed by large LXeTPCs experiments searching for dark matter particles in deep underground laboratories, such as XENON1T [59], PandaX-4T [31] and LUX [60]. The source of this background is not well understood yet, but we can roughly divide the delayed electrons into two groups.
The first kind of delayed electrons is produced via photon-ionization of impurities or metals in the drift region of the TPC. Photon-ionized delay electrons only appear shortly (within a time range that is comparable to the largest drift time of the LXeTPC) after the high energy S2 signals and will be completely subdominant when a S exposure is rejected for each of the high energy S2s, leading to an exposure loss of 1%.
Besides photon-ionization, delayed electrons can also be produced in a larger time range of up to a few seconds. The rate of delayed electrons is observed to be correlated with the level of impurities in LXe and the amplitude of the extraction field. Muons passing through the LXeTPC result in large energy deposition, generating a substantial number of ionization electrons. These delayed electrons may occur after the primary S2 signal of muon events, leading to a prolonged tail of signals composed of a single or a few electrons. Due to the high rate of cosmic muons, RELICS faces a major challenge in suppressing background from accidental pile-ups of delayed electrons in a short time range that is comparable to the width of a physical S2 signal. This background is hereafter referred to as Delayed Electron (DE).
V.5.1 Expected rate of DE
As shown in previous studies with XENON1T [59] and LUX [60], the emission rate of delayed electrons are correlated, in time and position, with the previous high-energy events. This part of the background is referred to as a correlated delayed electron background and can be significantly suppressed due to its unique correlation with proceeding high-energy events. We simulate muons passing through the TPC active volume using Geant4, tracking their interaction vertex and the energy deposited along the track. The ionization signals are generated along the muon track, assuming a constant ionization yield of 50 electrons per keV. We then simulated the generation of delayed electrons, assuming that 0.1% of the total ionization electrons from primary interactions are delayed emissions in a time range of 0.1 ms to 2 s, comparable to [59]. The distribution in “delayed time” (time respected to the primary S2) for delayed electrons is assumed to follow a power law with a coefficient of -1.1, as shown in XENON1T [59]. These delayed electrons’ horizontal positions () are assumed to be the same as their primary S2s. Position reconstruction resolution for single electron S2s, which is expected to be 20 mm from optical simulations, is taken into account in this analysis. Ionization S2 signals are grouped together if the time gap between them is below 1 s in order not to split S2 waveforms of events happening at the bottom part of the TPC. As a result, the pile-up of DE (called Pile-up DEs) leads to a significant background in CENS search. In RELICS, we expect more than 10 Hz of muon events in the LXeTPC, leaving a large amount of single electron S2 signals. The expected event rate of a single-electron pile-up background is summarized in Table 2. Without implementing appropriate cuts to reduce the rate of these pile-up events, CENS signals with 3 to 6 electrons will be completely overwhelmed.
| Ne | CENS | Pile-up DEs |
| 3 | 3.5 | 9.8 |
| 4 | 1.5 | 7.3 |
| 5 | 6.9 | 6.9 |
| 6 | 3.2 | 7.5 |
V.5.2 Waveform classifier
The waveform characteristics of Pile-up DE events differ from the physical interactions in the LXeTPC, such as CENS and other backgrounds. In physical interactions, electrons drift upwards under the drifting electric field, resulting in signal waveform distributions that follow a Gaussian distribution in time. As shown in figure 10, Pile-up DEs are distinguishable from physical interactions in terms of the shape of waveforms. The discrimination between them is maximized by employing a convolutional neural network (CNN), as shown in figure 11. While maintaining 80% of the physical interactions with a score over 0.8, the Pile-up DE events are reduced by approximately one order of magnitude.
V.5.3 Pattern classifier
Beyond the distinctions in the waveform or time distribution, the planar distribution of signals within the top PMT arrays also presents notable differences between Pile-up DE and physical interactions. CENS and other physical signals are composed of multiple electrons generated from the same position above the liquid surface in the detector, while the DE signals are mostly generated around the previous muons’ tracks. These DEs are merged together into larger signals only due to their temporal proximity. Therefore, the distributions of photo-electrons from CENS signals and DE signals differ in the top PMT array of the detector. Real physical interactions appear more like a superposition of multiple single-electron signals from the same location, whereas the Pile-up DEs are formed by the superposition of single-electron signals originating from multiple different horizontal positions, thus appearing more dispersed.
After each S2 signal in the detector, the expected photons to be detected in each channel on the top PMT array are denoted as . is determined by the total area of S2, the relative LCE in that channel, and the PMT responses including detection efficiency. LCE of each PMT is derived with light propagation simulation in the RELICS TPC. It is assumed that the number of photo-electrons received by the PMT channel follows a Poisson distribution with as the mean. The detected photo-electrons are further derived considering the resolution of PMTs to single photo-election.
Given the expected and observed light distribution pattern, we calculated the goodness-of-fit likelihood to determine whether the event is from physical interactions or Pile-up DEs. The logarithm of goodness-of-fit likelihood is written as:
where and are expected and observed photo-electrons in a specific channel.
The S2 distribution pattern on top PMT array of the Pile-up DEs rejected by a selection in (figure 12, left) would look like that of physical interactions (figure 12, right). Since both of horizontal position and emission time of DEs are strongly correlated to the previous muon tracks, we further calculate the “Space-time Correlation Coefficient” of each event with previous muons to distinguish Pileup DEs from physical interactions that are independent of any muon events in the RELICS LXeTPC.
For each energy deposition of muon passing through the RELICS TPC, the reconstructed position of DEs would follow a Gaussian distribution determined by the resolution of position reconstruction. The distribution of DEs in time with respect to the previous muon events is assumed to follow a power law, as seen in [59]. Assuming is the expected number of DE in a specific position, the probability distribution function of DEs can be expressed as:
where is the position of muon track, is the observed position, is the time delay to the muon track and is the resolution of reconstructed position. Considering that the emission of DE can last up to 2s and the muon event rate of 10 Hz, we sum up the contributions of from the previous 20 muons and define the log of this value as the “Space-time Coefficient”, which is defined as:
We simulated 3.1 muon events, corresponding to 3.5 days of live time, and calculated the Pattern Likelihood and “Space-time Coefficient” . The two-dimensional distribution for Pile-up DE and physical interactions, within an S2 range of 120 to 240 PE, are shown in figure 13.
Considering the correlation between , P and the S2 area of each event, we provide a score for each event by Eq.6. Events with higher scores are inclined to be from physical interactions.
While restricting the acceptance of Pile-up DE events lower than %, we scanned the parameter space of and , in order to retain as high signal acceptance as possible. The fitted values are of 9.10 , of 0.57 and of -65.78.
Figure 14 delineates the score distribution for Pile-up DEs and physical interactions. The final classifier identifies all events with scores exceeding 0 as real physical interactions. In the ROI of S2 ranging from 120 to 240 PE, the acceptance for physical interactions such as CENS is 67%. This cut reduced the rate of Pile-up DEs by 4 orders of magnitude in the same ROI, leaving the rate of Pile-up DE background below the expected CENS signal in the whole ROI.
V.6 Summary
The background distribution from CRNs, cosmic muons, and detector materials are converted to S2 size from recoil energy using RelicsAPT which is described in Sec. III. All the background components mentioned above are summarized in figure 15. The CENS event rate is overlaid. The dominant background contribution comes from the Pile-up DEs, despite its large discrimination power to physical events as described in V.5. The NR background contribution from CRNs is an order of magnitude higher than muons, which is the second major contribution. The detector materials contribute a relatively low and flat component in ER. The event rate of all the background components in 32 kgyr exposure is summarized in Tab. 3. A signal-to-background ratio greater than 6 can be achieved in [120, 240] PEs ROI of CENS detection.
| Event Type | Nevts |
| Pile-up DE | 1429.0 |
| CRNs | 470.4 |
| Muon-induced neutrons | 2.3 |
| ER | 232.0 |
| Total Background | 2133.7 |
| CENS | 13673.5 |
VI Physics potential
To explore the physics potential of the RELICS detector, this section will discuss its sensitivity to the weak mixing angle and non-standard interactions (NSI) based on the background estimation in section V with 32 kgyr exposure. Several systematic uncertainties are considered, including 5% reactor neutrino flux, 5% ER and NR background uncertainty. The uncertainty of the Qy introduced in Sec. III.1 is the dominant one in our estimation.
| Nuisance Parameters | Constraints | |
| Charge yield morpher | 1 [37] | |
| Reactor neutrino flux multiplier | ||
| ER background rate multiplier | ||
| Neutron background rate multiplier | ||
| Pile-up DE background rate multiplier | ||
VI.1 Weak mixing angle
A precise measurement of CENS cross-section allows for a constraint on the weak mixing angle . Also known as the Weinberg angle, is conventionally used in the form of in the comparison between theory and experiment. Since is given by , can be extracted from the measured cross-section, and any deviation from the SM prediction will indicate new physics.
The of likelihood for measurement is defined as Eq. 7 where index indicates different S2 intervals. and are the observed and expected signal events in each S2 interval, respectively. is the background rate estimation and index runs over background components, including ER and neutrons. The rest are nuisance parameters. and are the multipliers of the ER and neutron backgrounds, respectively. denotes the morphers of charge yields. A Gaussian constraint term is imposed on each nuisance parameter, as summarized in Tab. 4.
The expected constraint on from RELICS is shown in figure 16. The great improvement from the red solid line to the blue dashed line indicates that the charge yield variation is a major contribution to the constraint uncertainty. The sensitivity is expected to be competitive and improve the Dresden-II’s result [61]. As illustrated in Fig. 17, RELICS will provide a measurement on with momentum transfer down to the MeV scale, which is a valuable complementary to the current measurements.
VI.2 Nonstandard neutrino interactions
The detection of CENS provides a new probe to study neutrino physics beyond the SM, such as the NSI of neutrinos. NSI has been extensively studied in the literature [64]. This work focuses on a model-independent approach in the neutrino-quark sector. The parameterization of the NSI contributions to the cross-section can be expressed as Eq. 8, which indicates that the cross-section strongly depends on the effective electron neutrino-up quark and -down quark interaction strength parameters of and .
Eq. 9 defines the of likelihood, where all the parameters are the same as Eq. 7, except and . All the related uncertainties of nuisance parameters are also taken in Tab. 4. Figure 18 shows the constraining on and of RELICS under the standard model assumption, where both of the parameters are zero. The latest result from COHERENT [65] is overlaid for comparison. A tighter constraint can be achieved for the RELICS experiment with higher statistics.
VII Summary
The RELICS experiment is a (100) kg LXe detector with a fiducial mass of 32 kg that aims to observe the process of reactor CENS. The experiment is planned to be deployed at Sanmen Nuclear Power Plant (3 GW) in Zhejiang province, China, with a distance of 25 m, an average neutrino flux of 1013cm-2s-1.
In this work, a preliminary design of the RELICS experiment was presented. A dedicated MC simulation has been performed to study the background contributions. The dominant backgrounds come from the Pile-up DEs. The second largest background, nuclear recoils induced by cosmic-ray neutrons, is suppressed via a passive water shield. An ionization-only analysis channel will be adopted for RELICS analysis to achieve a sufficiently low energy threshold for CENS detection. A signal-to-background ratio of greater than 6 can be achieved in [120, 240] PEs ROI of CENS detection. Its sensitivity to weak mixing angle and NSI are explored with 32 kgyr exposure.
The RELICS experiment will provide rich opportunities to study Beyond the Standard Model (BSM) physics and potential applications in nuclear reactor monitoring for nuclear safeguards.
Acknowledgements. We acknowledge Sanmen Nuclear Power Plant for hosting RELICS. We thank Dr. Xun Chen for his support on the BambooMC package and Prof. Kaixuan Ni for fruitful discussions. RELICS is supported by a National Key R&D grant from the Ministry of Science and Technology of China (No. 2021YFA1601600) and a grant from Natural Science Foundation of China (No. 12275267). Yifei Zhao’s work on RELICS is supported by a grant for undergraduate research from Beijing Natural Science Foundation (No. QY23088).